OpenStudy (anonymous):

y is directly proportional to the square of x. When x = 3, y = 36.Find the value of y when x = 5

OpenStudy (anonymous):

help

OpenStudy (ddcamp):

If y is directly proportional to the square of x, there is some number you multiply x² by to get x: y=kx² This constant, k, will be the same for any value of x. Can you find k if you know that y=36 when x=3? Hint: Use the known values for x and y in the formula y=kx²

OpenStudy (anonymous):

y=kx² 36=k3² 36=9k k=36/9 k=4k y=4²k???

OpenStudy (anonymous):

is that right so far

OpenStudy (anonymous):

y is directly proportional to the square of x. When x = 3, y = 36.Find the value of y when x = 5 help

OpenStudy (anonymous):

help please

OpenStudy (ddcamp):

So, you found that k=4. Well done! Now, just plug that into the equation, but with x=5: y=kx² y=4(5²) y=4(25) y=100

OpenStudy (anonymous):

thank you..that is where I was getting confused

OpenStudy (anonymous):

hw do you give someone a medal

OpenStudy (anonymous):

 A, B, C and D are points on the circumference of a circle, centre O. Angle AOC = y. Find the size of angle ABC in terms of y. Give a reason for each stage of your working.